OpenStudy (anonymous):

HELP PLEASE!!!!!! Th angle of elevation to the top of a building is 21 degrees. At a point 80 feet closer, the angle of elevation to the top of the same building is 33 degrees. Approximate the height of the building.

OpenStudy (anonymous):

\[since:\tan 33 = h/x \] \[\therefore h = x \tan 33\] --->1 \[since \tan 21 = h / (80+x)\] \[\therefore h = (80+x)\tan 21\]---->2 \[solve 1 & 2 together\] \[xtan33=80\tan21+xtan21\] \[-80\tan21 = -xtan33+xtan21\] \[80 \tan21= x(\tan33-\tan21)\] \[x=115.6462729 substitue In ---->1\] \[h= 75.10156777 feets \approx 75.1 feets\]